# Flex-Judge: Think Once, Judge Anywhere

### v1

@@ 4 Broader Impact: Case Study on Molecule Evaluator
A particularly compelling use case arises in scientific domains such as molecular modeling [3, 56, 79], where no existing reward model or judge is available due to limited data and domain complexity. To test our hypothesis, we build a molecular judge using Mol-LLaMA [27], which is based on LLaMA3.1-8B [23] and understands both 2D and 3D molecule representations. We fine-tune its LLM backbone as a reasoning-aware judge, while keeping its modality-aware modules (LoRA adapters [26], molecular encoders, and Q-Formers [37]) unchanged. We refer to this model as Flex-Mol-LLaMA.
We evaluate Flex-Mol-LLaMA via two approaches. (1) Best-of-$N$ selector: Flex-Mol-LLaMA guides best-of-$N$ sampling from Mol-LLaMA outputs, a method correlated with judge model performance [38]. (2) Direct preference optimization (DPO): Given a prompt $\mathbf{x}$, Mol-LLaMA generates two responses $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$, and Flex-Mol-LLaMA judge selects the preferred ($\mathbf{y}_{w}$) over the less preferred ($\mathbf{y}_{l}$) by score evaluation, forming triplets $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$ [50]. These are then used to fine-tune Mol-LLaMA via DPO [54]. Strong downstream performance indicates that Flex-Mol-LLaMA effectively curates high-quality preference data.
Best-of-$N$ Selector.
We use Flex-Mol-LLaMA to score the Mol-LLaMA’s understanding of chemical property (permeability), following Kim et al. [27]. Our key findings include:
@@
Mol-Instructions [22] 55.91 33.50 70.47
3D-MoLM [39] 46.93 50.00 64.86
LLaMo [51] 49.25 64.37 48.51
Mol-LLaMA [27] 63.55 64.37 72.48
@@
Figure 4: (Left) Accuracy (%) trends on the parallel artificial membrane permeability assay (PAMPA; [61]) task with different judgment scores. (Middle) Accuracy trends on the number of sampled responses in best-of-$N$ sampling. (Right) Performance comparison with reward-guided Mol-LLaMA. We report accuracy with prompt styles of default, CoT, and task information, as described in [27].
@@
  • Best-of-$N$ Sampling: Selecting the highest-scoring ones among $N$ sampled responses yields a significant accuracy improvement (Figure 4; middle, up to 77.49% when $N=16$), indicating that our judge model provides reliable preference signals.
@@ 5 Analysis
To investigate the impact of reasoning in Flex-Judge, we compare it against a non-reasoning variant where the training segments <think> and <answer> are reversed—that is, the model is trained to generate the final answer first, followed by the reasoning. This answer-then-reason format is commonly used in open-source datasets [86, 73]. As shown in Figure 5 (left), our proposed Flex-Judge consistently outperforms this variant, demonstrating that reasoning-first evaluation leads to more robust and generalizable performance. Furthermore, in Figure 5 (right), we find that accuracy gains on MLLM-as-a-Judge correlate with the average reasoning length produced by Flex-VL-7B. This suggests that more difficult tasks elicit deeper reasoning, highlighting the importance of reasoning capabilities in accurate evaluation. Notably, the M2W dataset [19], which involves complex website browsing scenarios and thus requires fine-grained reasoning paths, stands out with a larger performance gain than the general trend.
@@
Figure 6: Inference-time scaling of Flex-(VL/Omni)-7B. Unlike prior MLLM evaluators [38], our model supports parallel inference-time scaling via majority voting, showing consistent gains. Surprisingly, budget forcing also offers minor but consistent improvements in score-based evaluations.
@@
We examine inference-time scaling techniques, including majority voting [66] and self-refinement through budget forcing [44, 49], to enhance the reasoning ability of our Flex-Judge. In the VL-RewardBench (Reasoning) pairwise comparison, applying majority voting steadily improves performance, which is in stark contrast to prior work [38] that reported performance drops under inference-time scaling (6(a)). For budget forcing, we inject the keyword "Wait", shown to be effective in Muennighoff et al. [49]. Although performance drops after the first trial, it consistently improves in subsequent trials and nearly recovers to its original level (6(b)). In score-based evaluation, both methods show consistent gains (6(c) and 6(d)). These results highlight that, unlike existing MLLM-based judges, our evaluator benefits from increased inference-time computation, especially in reasoning-heavy tasks. Thanks to its reasoning-based training, our Flex-Judge produces more diverse reasoning paths than prior non-reasoning evaluators, allowing inference-time scaling methods to be more effective.
@@
To further assess the effectiveness of our cost-efficient approach, we compare Flex-VL-7B with a variant trained on image-text evaluation pairs. Using the RLHF-V dataset [81], we curated 1K image-text pairs with reasoning-guided evaluation where Flex-VL-7B judged the chosen response as better—without relying on explicit GPT-4o-annotated scores, potentially resulting in weaker quality than our original dataset (). As shown in Table 5, these variants consistently underperform Flex-Judge, emphasizing that dataset quality outweighs modality-awareness.
@@ 6 Related Work
LLMs have increasingly been used as proxy evaluators in place of costly human annotators, a direction formalized under the LLM-as-a-Judge paradigm [85, 8], which has demonstrated strong alignment with human preferences [58, 57, 60, 28]. These model-based judgments can be leveraged in various downstream techniques such as best-of-$N$ selection [24] and preference-driven optimization using DPO [54, 62] or listwise reranking [43, 53]. However, most LLM-as-a-Judge approaches remain limited to text-only domains [36], while their multimodal variants (e.g., ImageReward [75]) require large-scale, modality-specific annotations [6, 73, 32]. Recent work has highlighted the advantages of reasoning-guided supervision, training models with explanations or chain-of-thought rationales [7, 69, 68], to improve judgment quality and generalization [12]. Yet, collecting high-quality multimodal rationales is costly and especially difficult for underexplored modalities [11, 30, 84], limiting applicability to new domains. Our work addresses this gap by showing that textual reasoning alone can effectively train multimodal judge models to generalize across modalities, including tasks like molecular evaluation where no modality-specific preference data exists [27]. Additional related works are discussed in Appendix A.
@@ Appendix A Additional Related Work
Large language models (LLMs) have recently been explored as a cost-effective alternative to human raters, a paradigm referred to as LLM-as-a-Judge [85]. By encoding human-like reasoning and leveraging instruction prompts, these models can approximate human preference judgments in tasks such as summarization, dialogue, or code generation [58, 57, 29, 60]. Because human annotations are expensive and time-consuming to scale, the LLM-based evaluation promises a reusable, modular approach that can reduce reliance on direct human supervision. For instance, BoNBoN Alignment [24] explores using best-of-N selection guided by LLM-generated scores, while LRL [43] relies on LLM-based ranking signals to optimize model outputs at inference time.
However, most LLM-as-a-Judge approaches have focused on text-only use cases [36], and extending these models to multimodal content (e.g., vision-language or speech tasks) remains challenging [10, 11, 65]. While multimodal variants do exist [75, 73], they often rely on large-scale, manually crafted, and modality-specific annotated datasets for training or fine-tuning [6, 84, 74], which becomes overly expensive at scale. For example, LLaVA-Critic [73] curates 46K images and 113K evaluation data by using GPT-4/4V on 8 multimodal datasets, as well as manually crafted prompt templates. Prometheus-Vision [32] involves a number of data processing steps, with 5K images and 15K customized score rubrics, and GPT-4V generating 150K response-feedback pairs.
Furthermore, the lack of high-quality, publicly available multimodal preference benchmarks makes it difficult to develop MLLM judge models [11, 30]. Our work addresses this gap by demonstrating that an MLLM can effectively judge multimodal outputs without requiring extensive modality-specific preference supervision. This approach suggests that LLMs as judge models can be extended to a wider range of modalities (e.g., 3D point clouds [25], LiDAR [77], molecular [27, 79]) that have not yet been explored due to limited resources.
@@ A.2 Reasoning-Guided Reward Models
A growing body of research suggests that equipping reward models with the ability to reason—often captured through chain-of-thought prompts or explicit textual rationales—can significantly improve alignment with human preferences [58, 7, 69]. Rather than learning only from binary or scalar labels, these models benefit from learning how humans arrive at a decision, providing a more interpretable and generalizable internal mechanism for preference modeling [55, 31].
Recent advances in this area have shown that training on textual explanations can improve zero-shot and few-shot performance in downstream tasks requiring nuanced judgments [57, 29]. For example, Zheng et al. [85] find that when LLMs articulate the rationale behind their preference for one sample over another, they achieve higher consistency with human annotators. Building on the success of reasoning-based models in other domains, Zhang et al. [83] and Mahan et al. [45] introduced generative reward models, which significantly outperform non-reasoning judge models. More recently, JudgeLRM [12] demonstrates that reinforcement learning with reasoning-annotated data leads to substantial gains on reasoning-intensive evaluation tasks, significantly outperforming standard SFT-trained models. Concurrent with JudgeLRM, models such as J1 [70], DeepSeek-GRM [42], and RM-R1 [13] also adopt reinforcement learning frameworks to enhance the reasoning capabilities of reward models.
Despite these benefits, most reasoning-based reward modeling remains constrained to text-only applications, due in part to the higher cost and complexity of collecting multimodal rationales [73, 11]. In contrast, our work shows that textual reasoning supervision alone can be leveraged to enable robust multimodal evaluation. By training models on a small set of high-quality textual explanation data, we demonstrate an effective cross-modal generalization without the need for domain-specific preference labels. This approach reduces annotation overhead and paves the way for more scalable, modality-agnostic judge models.
@@ B.1 Dataset Description
Figure 7: System prompt for JudgeLRM [12].
We first describe our training dataset (seed): To generate reasoning-based judgments, we use JudgeLRM-7B [12] to sample responses based on the given contexts from JudgeLM-100K [86], using the prompts shown in Figure 7 with a temperature of $0.1$. Among the 100K samples, we filter out those with rating mismatches compared to the GPT-4o evaluation results provided in Zhu et al. [86], resulting in approximately 20K samples remaining after this process.
@@
  • MLLM-as-a-Judge: The MLLM-as-a-Judge benchmark [11] is introduced to specifically assess the judgment ability of MLLMs in the domain of image understanding across diverse scenarios. The benchmark consists of 14 datasets covering tasks such as image captioning, mathematical reasoning, text recognition, and infographic understanding. In total, it includes 4,414 image-instruction pairs, curated to evaluate whether MLLMs can generalize their evaluative abilities across different modalities.
  • VL-RewardBench: VL-RewardBench [38] is a diagnostic benchmark for evaluating vision-language models on multimodal understanding, hallucination detection, and complex reasoning. It comprises 1,250 high-quality examples curated through an AI-assisted pipeline with human verification.
  • MJ-Bench: MJ-Bench [14] is a benchmark designed to evaluate multimodal foundation models in the role of a judge for image generation tasks. It includes preference data across four key perspectives—text-image alignment, safety, image quality, and generation bias—each further divided into detailed subcategories. Each example consists of an instruction paired with a chosen and rejected image, enabling fine-grained assessment of model judgment.
  • GenAI-Bench: GenAI-Bench [35] is a human-curated benchmark for evaluating image and video generation models across diverse composition skills. It includes 1,600 prompts from professional designers, avoiding subjective or inappropriate content, and covers over 5,000 human-verified skill tags. Unlike prior work, each prompt is annotated with multiple fine-grained tags. Specifically, GenAI-Bench supports image generation, image editing, and video generation tasks.
  • Audio MOS/SS Benchmark: There is no unified, structured benchmark for audio evaluation tasks. Instead, Wang et al. [65] assessed speech quality and speaker similarity using four datasets: NISQA [48], BVCC [18], and SOMOS [46] for speech quality (712, 742, and 3,000 test samples, respectively), and VoxSim [2] for speaker similarity (2,776 test pairs). All datasets include human-annotated scores. For the MOS prediction task, auditory LLMs are asked to assign the MOS score on a scale from 1.0 to 5.0 for a given speech input. While Wang et al. [65] designed dataset-specific prompts to help models account for each dataset’s unique standards, we evaluate models using a unified prompt format without dataset-specific tuning. For the SS prediction task, models rate the similarity between two speech samples on a scale from 1.0 to 6.0, where higher scores indicate greater speaker similarity. Since the human annotations of MOS and SS are averaged over multiple individuals, they are predicted in the form of floats.

### v3

@@ 4 Broader Impact: Case Study on Molecule Evaluator
A particularly compelling use case arises in scientific domains such as molecular modeling AI4Science and Quantum (2023); Sadeghi et al. (2024); Yu et al. (2024a), where no existing reward model or judge is available due to limited data and domain complexity. To test our hypothesis, we build a molecular judge using Mol-LLaMA Kim et al. (2025), which is based on LLaMA3.1-8B Grattafiori et al. (2024) and understands both 2D and 3D molecule representations. We fine-tune its LLM backbone as a reasoning-aware judge, while keeping its modality-aware modules (LoRA adapters Hu et al. (2022), molecular encoders, and Q-Formers Li et al. (2023)) unchanged. We refer to this model as Flex-Mol-LLaMA.
We evaluate Flex-Mol-LLaMA via two approaches. (1) Best-of-$N$ selector: Flex-Mol-LLaMA guides best-of-$N$ sampling from Mol-LLaMA outputs, a method correlated with judge model performance Li et al. (2024c). (2) Direct preference optimization (DPO): Given a prompt $\mathbf{x}$, Mol-LLaMA generates two responses $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$, and Flex-Mol-LLaMA judge selects the preferred ($\mathbf{y}_{w}$) over the less preferred ($\mathbf{y}_{l}$) by score evaluation, forming triplets $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$ Ouyang et al. (2022). These are then used to fine-tune Mol-LLaMA via DPO Rafailov et al. (2023). Strong downstream performance indicates that Flex-Mol-LLaMA effectively curates high-quality preference data.
Best-of-NN Selector.
We use Flex-Mol-LLaMA to score the Mol-LLaMA’s understanding of chemical property (permeability), following Kim et al. (2025). Our key findings include:
@@
Mol-Instructions Fang et al. (2024) 55.91 33.50 70.47
3D-MoLM Li et al. (2024d) 46.93 50.00 64.86
LLaMo Park et al. (2024) 49.25 64.37 48.51
Mol-LLaMA Kim et al. (2025) 63.55 64.37 72.48
@@
Figure 4: (Left) Accuracy (%) trends on the parallel artificial membrane permeability assay (PAMPA; Velez-Arce et al. (2024)) task with different judgment scores. (Middle) Accuracy trends on the number of sampled responses in best-of-NN sampling. (Right) Performance comparison with reward-guided Mol-LLaMA. We report accuracy with prompt styles of default, CoT, and task information, as described in Kim et al. (2025).
@@
  • Best-of-NN Sampling: Selecting the highest-scoring ones among $N$ sampled responses yields a significant accuracy improvement (Figure 4; middle, up to 77.49% when $N=16$), indicating that our judge model provides reliable preference signals.
@@ 5 Analysis
To investigate the impact of reasoning in Flex-Judge, we compare it against a non-reasoning variant where the training segments <think> and <answer> are reversed—that is, the model is trained to generate the final answer first, followed by the reasoning. This answer-then-reason format is commonly used in open-source datasets Zhu et al. (2025); Xiong et al. (2024). As shown in Figure 5 (left), our proposed Flex-Judge consistently outperforms this variant, demonstrating that reasoning-first evaluation leads to more robust and generalizable performance. Furthermore, in Figure 5 (right), we find that accuracy gains on MLLM-as-a-Judge correlate with the average reasoning length produced by Flex-VL-7B. This suggests that more difficult tasks elicit deeper reasoning, highlighting the importance of reasoning capabilities in accurate evaluation. Notably, the M2W dataset Deng et al. (2023), which involves complex website browsing scenarios and thus requires fine-grained reasoning paths, stands out with a larger performance gain than the general trend.
@@
Figure 6: Inference-time scaling of Flex-(VL/Omni)-7B. Unlike prior MLLM evaluators Li et al. (2024c), our model supports parallel inference-time scaling via majority voting, showing consistent gains. Surprisingly, budget forcing also offers minor but consistent improvements in score-based evaluations.
@@
We examine inference-time scaling techniques, including majority voting Wang et al. (2023b) and self-refinement through budget forcing Madaan et al. (2023); Muennighoff et al. (2025), to enhance the reasoning ability of our Flex-Judge. In the VL-RewardBench (Reasoning) pairwise comparison, applying majority voting steadily improves performance, which is in stark contrast to prior work Li et al. (2024c) that reported performance drops under inference-time scaling (6(a)). For budget forcing, we inject the keyword "Wait", shown to be effective in Muennighoff et al. (2025). Although performance drops after the first trial, it consistently improves in subsequent trials and nearly recovers to its original level (6(b)). In score-based evaluation, both methods show consistent gains (6(c) and 6(d)). These results highlight that, unlike existing MLLM-based judges, our evaluator benefits from increased inference-time computation, especially in reasoning-heavy tasks. Thanks to its reasoning-based training, our Flex-Judge produces more diverse reasoning paths than prior non-reasoning evaluators, allowing inference-time scaling methods to be more effective.
@@
To further assess the effectiveness of our cost-efficient approach, we compare Flex-VL-7B with a variant trained on image-text evaluation pairs. Using the RLHF-V dataset Yu et al. (2024b), we curated 1K image-text pairs with reasoning-guided evaluation where Flex-VL-7B judged the chosen response as better—without relying on explicit GPT-4o-annotated scores, potentially resulting in weaker quality than our original dataset (▲). As shown in Table 5, these variants consistently underperform Flex-Judge, emphasizing that dataset quality outweighs modality-awareness.
@@ 6 Related Work
LLMs have increasingly been used as proxy evaluators in place of costly human annotators, a direction formalized under the LLM-as-a-Judge paradigm Zheng et al. (2023); Bavaresco et al. (2024), which has demonstrated strong alignment with human preferences Stiennon et al. (2020); Song et al. (2025); Tong and Zhang (2024); Kim et al. (2023). These model-based judgments can be leveraged in various downstream techniques such as best-of-$N$ selection Gui et al. (2024) and preference-driven optimization using DPO Rafailov et al. (2023); Wallace et al. (2024) or listwise reranking Ma et al. (2023); Qin et al. (2024). However, most LLM-as-a-Judge approaches remain limited to text-only domains Li et al. (2024b), while their multimodal variants (e.g., ImageReward Xu et al. (2023)) require large-scale, modality-specific annotations Bai et al. (2024); Xiong et al. (2024); Lee et al. (2024). Recent work has highlighted the advantages of reasoning-guided supervision, training models with explanations or chain-of-thought rationales Bai et al. (2022); Wei et al. (2022); Wang et al. (2023c), to improve judgment quality and generalization Chen et al. (2025b). Yet, collecting high-quality multimodal rationales is costly and especially difficult for underexplored modalities Chen et al. (2024a); Kirstain et al. (2023); Zhang et al. (2025), limiting applicability to new domains. Our work addresses this gap by showing that textual reasoning alone can effectively train multimodal judge models to generalize across modalities, including tasks like molecular evaluation where no modality-specific preference data exists Kim et al. (2025). Additional related works are discussed in Appendix A.
@@ Appendix A Additional Related Work
Large language models (LLMs) have recently been explored as a cost-effective alternative to human raters, a paradigm referred to as LLM-as-a-Judge Zheng et al. [2023]. By encoding human-like reasoning and leveraging instruction prompts, these models can approximate human preference judgments in tasks such as summarization, dialogue, or code generation Stiennon et al. [2020], Song et al. [2025], Kim et al. [2024], Tong and Zhang [2024]. Because human annotations are expensive and time-consuming to scale, the LLM-based evaluation promises a reusable, modular approach that can reduce reliance on direct human supervision. For instance, BoNBoN Alignment Gui et al. [2024] explores using best-of-N selection guided by LLM-generated scores, while LRL Ma et al. [2023] relies on LLM-based ranking signals to optimize model outputs at inference time.
However, most LLM-as-a-Judge approaches have focused on text-only use cases Li et al. [2024b], and extending these models to multimodal content (e.g., vision-language or speech tasks) remains challenging Chen et al. [2025a, 2024a], Wang et al. [2025]. While multimodal variants do exist Xu et al. [2023], Xiong et al. [2024], they often rely on large-scale, manually crafted, and modality-specific annotated datasets for training or fine-tuning Bai et al. [2024], Zhang et al. [2025], Xu et al. [2024], which becomes overly expensive at scale. For example, LLaVA-Critic Xiong et al. [2024] curates 46K images and 113K evaluation data by using GPT-4/4V on 8 multimodal datasets, as well as manually crafted prompt templates. Prometheus-Vision Lee et al. [2024] involves a number of data processing steps, with 5K images and 15K customized score rubrics, and GPT-4V generating 150K response-feedback pairs.
Furthermore, the lack of high-quality, publicly available multimodal preference benchmarks makes it difficult to develop MLLM judge models Chen et al. [2024a], Kirstain et al. [2023]. Our work addresses this gap by demonstrating that an MLLM can effectively judge multimodal outputs without requiring extensive modality-specific preference supervision. This approach suggests that LLMs as judge models can be extended to a wider range of modalities (e.g., 3D point clouds Hong et al. [2023], LiDAR Yang et al. [2025], molecular Kim et al. [2025], Yu et al. [2024a]) that have not yet been explored due to limited resources.
@@ A.2 Reasoning-Guided Reward Models
A growing body of research suggests that equipping reward models with the ability to reason—often captured through chain-of-thought prompts or explicit textual rationales—can significantly improve alignment with human preferences Stiennon et al. [2020], Bai et al. [2022], Wei et al. [2022]. Rather than learning only from binary or scalar labels, these models benefit from learning how humans arrive at a decision, providing a more interpretable and generalizable internal mechanism for preference modeling Rajani et al. [2019], Lampinen et al. [2022].
Recent advances in this area have shown that training on textual explanations can improve zero-shot and few-shot performance in downstream tasks requiring nuanced judgments Song et al. [2025], Kim et al. [2024]. For example, Zheng et al. [2023] find that when LLMs articulate the rationale behind their preference for one sample over another, they achieve higher consistency with human annotators. Building on the success of reasoning-based models in other domains, Zhang et al. [2024] and Mahan et al. [2024] introduced generative reward models, which significantly outperform non-reasoning judge models. More recently, JudgeLRM Chen et al. [2025b] demonstrates that reinforcement learning with reasoning-annotated data leads to substantial gains on reasoning-intensive evaluation tasks, significantly outperforming standard SFT-trained models. Concurrent with JudgeLRM, models such as J1 Whitehouse et al. [2025], DeepSeek-GRM Liu et al. [2025], and RM-R1 Chen et al. [2025c] also adopt reinforcement learning frameworks to enhance the reasoning capabilities of reward models.
Despite these benefits, most reasoning-based reward modeling remains constrained to text-only applications, due in part to the higher cost and complexity of collecting multimodal rationales Xiong et al. [2024], Chen et al. [2024a]. In contrast, our work shows that textual reasoning supervision alone can be leveraged to enable robust multimodal evaluation. By training models on a small set of high-quality textual explanation data, we demonstrate an effective cross-modal generalization without the need for domain-specific preference labels. This approach reduces annotation overhead and paves the way for more scalable, modality-agnostic judge models.
@@ B.1 Dataset Description
Figure 7: System prompt for JudgeLRM Chen et al. [2025b].
We first describe our training dataset (seed): To generate reasoning-based judgments, we use JudgeLRM-7B Chen et al. [2025b] to sample responses based on the given contexts from JudgeLM-100K Zhu et al. [2025], using the prompts shown in Figure 7 with a temperature of $0.1$. Among the 100K samples, we filter out those with rating mismatches compared to the GPT-4o evaluation results provided in Zhu et al. [2025], resulting in approximately 20K samples remaining after this process.
@@
  • MLLM-as-a-Judge: The MLLM-as-a-Judge benchmark Chen et al. [2024a] is introduced to specifically assess the judgment ability of MLLMs in the domain of image understanding across diverse scenarios. The benchmark consists of 14 datasets covering tasks such as image captioning, mathematical reasoning, text recognition, and infographic understanding. In total, it includes 4,414 image-instruction pairs, curated to evaluate whether MLLMs can generalize their evaluative abilities across different modalities.
  • VL-RewardBench: VL-RewardBench Li et al. [2024c] is a diagnostic benchmark for evaluating vision-language models on multimodal understanding, hallucination detection, and complex reasoning. It comprises 1,250 high-quality examples curated through an AI-assisted pipeline with human verification.
  • MJ-Bench: MJ-Bench Chen et al. [2024b] is a benchmark designed to evaluate multimodal foundation models in the role of a judge for image generation tasks. It includes preference data across four key perspectives—text-image alignment, safety, image quality, and generation bias—each further divided into detailed subcategories. Each example consists of an instruction paired with a chosen and rejected image, enabling fine-grained assessment of model judgment.
  • GenAI-Bench: GenAI-Bench Li et al. [2024a] is a human-curated benchmark for evaluating image and video generation models across diverse composition skills. It includes 1,600 prompts from professional designers, avoiding subjective or inappropriate content, and covers over 5,000 human-verified skill tags. Unlike prior work, each prompt is annotated with multiple fine-grained tags. Specifically, GenAI-Bench supports image generation, image editing, and video generation tasks.
  • Audio MOS/SS Benchmark: There is no unified, structured benchmark for audio evaluation tasks. Instead, Wang et al. [2025] assessed speech quality and speaker similarity using four datasets: NISQA Mittag et al. [2021], BVCC Cooper and Yamagishi [2021], and SOMOS Maniati et al. [2022] for speech quality (712, 742, and 3,000 test samples, respectively), and VoxSim Ahn et al. [2024] for speaker similarity (2,776 test pairs). All datasets include human-annotated scores. For the MOS prediction task, auditory LLMs are asked to assign the MOS score on a scale from 1.0 to 5.0 for a given speech input. While Wang et al. [2025] designed dataset-specific prompts to help models account for each dataset’s unique standards, we evaluate models using a unified prompt format without dataset-specific tuning. For the SS prediction task, models rate the similarity between two speech samples on a scale from 1.0 to 6.0, where higher scores indicate greater speaker similarity. Since the human annotations of MOS and SS are averaged over multiple individuals, they are predicted in the form of floats.